OpenStudy (anonymous):

Need help confusing problem Solve trig Equation: 2cosxcos2x+sinxsin2x+cosx=0 I can seem to break it down and use the correct identities)

OpenStudy (tkhunny):

What have you tried? I thought this was cute: 2cosxcos2x+sinxsin2x+cosx=0 cosxcos2x+cosxcos2x+sinxsin2x+cosx=0 cosxcos2x+(cosxcos2x+sinxsin2x)+cosx=0 cosxcos2x+(cos(x-2x))+cosx=0 cosxcos2x+cos(-x)+cosx=0 cosxcos2x+cos(x)+cosx=0 cosxcos2x+2cos(x)=0 cosx(cos2x+2)=0 Of course, if you really meant \(\cos^{2}(x)\) and not \(\cos(2x)\) then that's another story.

OpenStudy (tkhunny):

It is amazing how much better one can respond to a question when the entire question is presented. It's even better when you show your work!